if we have a pulley from which 2 weights of 2kg each are suspended at the right hand side and a weight of 2kg is suspended at the left side. what is the tension in the string between two masses at the right side?

Grade:12

2 Answers

Rohith Gandhi
14 years ago

Dear Kriti,

Consider the pulley and two masses of 2 kg each suspended to right and a single mass of 2 kg is suspended on the left. let tensions and accelarations be as shown in the figure.

Your question was as per figure what is the tension betwwen two blocks on the right which is nothing but T2. Applying newtons laws i will be able to solve the question.

5450-1036_8202_pulley.bmp

blocks on the right : top 2 kg block writing newtons second law,

2g-T1+T2= 2a -------eqn 1

2 kg block on the right below the top one,

2g - T2-= 2a ---------eqn 2

2kg block on the left,

T1-2g=2a-----------eqn3

substiting for T1 from eqn3 in eqn1 i.e. T1 = 2g +2a

=>2g - 2g-2a + T2 = 2a

=>T2=4a=>a=T2/4

substituting the above obtained a in eqn 2

=>2g-T2=2XT2/4

=>2g =1.5 T2

=> T2=4g/3

hence the tension between the two 2 kg blocks on the right side is 4g/3

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Kriti!!!




Regards,

Askiitians Experts
Rohtih Gandhi P

 

Askiitian.Expert Rajat
14 years ago
Hi Kriti, The tension in the string between the two masses will be equal to the weight of the lower mass that is it will be equal to mg where m=2.5kg and g = 9.8 m/s2 the force will be in newtons which is the si unit of force. Regards Rajat Askiitians.com Expert